Joey has two investmentstotaling $15000 he receives 8%from one and 7% iterest from the other. The 8% investment yields an anual income that is $150 less than the income from the 7% investmentfind the amount invested from each rate

Answer by richwmiller(17219)

x+y=15000,
0.08*x-0.07*y=150
x=15000-y
0.08*(15000-y)-0.07*y=150
1200-0.08y-0.07*y=150
-0.15*y=-1050
y=7000 at -7%
x=15000-y
x=8000 at 8%
check
0.08*8000-0.07*7000=150
640-490=150
150=150
ok